Story Problem: My initial investment is $4000.00 and is componded quarterly at 6.25% . How long will it take to double my money ?
Answer by stanbon(75887)

Story Problem: My initial investment is $4000.00 and is componded quarterly at 6.25% . How long will it take to double my money ?
---------------
2*4000= 4000(1+(0.0625/4)(4t)
2 = (1+(0.0625/4)(4t)
2 = (1.015625)^(4t)
Take the natural log to get:
4t = ln2/ln1.015626
4t = 44.70426
t=11.18 years
===============
Cheers,
Stan H.